OpenStudy (anonymous):

Identify a and b (as used in the above sum and difference of cubes expressions) in each of the following (do not factorise). a) 8y^3 + z^3 b) 1 − 64h^3 c) 1/8(c^3) - e^3g^3 d) t^3 - u^3/216 Show all working out thanks.

OpenStudy (anonymous):

a = 2y, b = z a = 1, b = 4h a = 1/(2c), b = eg a = t, b = u/6

OpenStudy (anonymous):

(2y)^3 + (z)^3 = 8y^3 + z^3 (1)^3 - (4h)^3 = 1 - 64h^3 ....